x+6/x^2+x-20
What value or values of x are not allowed in this rational expression?
(If you find two solutions, separate them with a comma.)

Found 2 solutions by ankor@dixie-net.com, solver91311:
Answer by ankor@dixie-net.com(22740)

x+6/x^2+x-20
What value or values of x are not allowed in this rational expression?
(If you find two solutions, separate them with a comma.)
:
Denominator cannot = 0, therefore
x^2 + x - 20 = 0
Factors to
(x+5)(x-4) = 0
x = -5
x = 4
-5, 4 values are not allowed in this expression

Answer by solver91311(24713)



Take the denominator function and set it equal to zero.

Solve the quadratic.

The two values obtained are the two values that make the rational expression undefined because if either value is substituted for x, the denominator of the rational function becomes zero.


John

My calculator said it, I believe it, that settles it